Many random number generators allow users to specify the range of the random numbers to be produced. Suppose that you specify that the random number Y can take any value between 0 and 2. Then the density curve of the outcomes has constant height between 0 and 2, and height 0 elsewhere.

B) What is the height of the density curve between 0 and 2? Draw a graph of the density curve.

C)Use your graph from (b) and the fact that probability is area under the curve to find P(Y ≤ 1).
Answer #1

b)as area under the density curve is =1

hence height *base =1

height *2=1

height =1/2 =0.5

c)

from above P(Y ≤ 1). =P(area under the curve below Y=1) =P(Area under the rectangle from 0 to 1)=height*base=0.5*(1-0)

=0.5